OpenStudy (anonymous):

10 years hence, A will be twice as old as B was 10 years ago. If A is now 10 years older than B, find the present age of B. I formed this equation for solving this sum: x + 9 + 10 = 2(x - 10) Where x + 9 is A's age and x is B's age. Just wanted to know if I am right.

OpenStudy (anonymous):

I got B's age as 39. Am I right??

OpenStudy (lalaly):

is this the full question?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

A+10=2(b-10)

OpenStudy (anonymous):

or actually, it's kind of confusing

OpenStudy (anonymous):

10 years from the future or the present?

OpenStudy (anonymous):

x + 9 + 10 denotes A's age 10 years hence. It is equal to 2(x - 10) which is twice of B's age 10 years ago.

OpenStudy (lalaly):

im confused too

OpenStudy (anonymous):

i guess ten years ago from 10 years in the future is the present, lol

OpenStudy (anonymous):

You have a point there.

OpenStudy (anonymous):

But then, they've given 10 years hence for A and 10 years ago for B.

OpenStudy (anonymous):

Pretty sure the original equation i gave is correct

OpenStudy (anonymous):

where did you get x + 9 from?

OpenStudy (anonymous):

Which is same as mine right? Only thing I've replaced A with x + 9.

OpenStudy (anonymous):

Well actually, they're not the same

OpenStudy (anonymous):

x + 9 is A's present age(B's present age is x and A is 9 years older than B) .

OpenStudy (anonymous):

I don't see anything about a 9 in the question...

OpenStudy (anonymous):

Oh, I made a typing mistake its actually if A is 9 years older than B, find B's present age.....

OpenStudy (anonymous):

Can you tell how? Because on solving the equation, I got 39 .....

OpenStudy (anonymous):

let me double check

OpenStudy (anonymous):

I'm sorry, you are right. B is presently 39 years old.

OpenStudy (anonymous):

Thanks for helping.

OpenStudy (anonymous):

A+10=2(B-10) A=2B-30 A=B+9 B+9=2B-30 B=39

OpenStudy (anonymous):

No problem